OpenStudy (anonymous):

Six liters of juice filled a container 2 1/2 times. How many liters of juice does the container hold?

OpenStudy (anonymous):

6* 2 1/2= ?

OpenStudy (anonymous):

15

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

I don't know if it's that simple.. is it?

OpenStudy (anonymous):

it is

OpenStudy (anonymous):

Ok I guess I was over-thinking it. thank you!

OpenStudy (anonymous):

ups nevr mind i made a mistake...

OpenStudy (apoorvk):

nope @myko if 6 litres filled the jar 2.5 times, then, the jar's capacity is 6/2.5. think it over.

OpenStudy (anonymous):

@apoorvk is right

OpenStudy (anonymous):

Would the answer be a fraction between 18 and 24 liters?

OpenStudy (anonymous):

6/5/2

OpenStudy (apoorvk):

6/2.5 is something like - 2.4

OpenStudy (anonymous):

6/5/2=12/5

OpenStudy (apoorvk):

Umm @bam56 that 4.5 isn't exactly double of 2.5, when you're doubling 6.

OpenStudy (anonymous):

Oh! I didn't catch that, I was going up by 4's. I'll rework on it.

OpenStudy (anonymous):

Ok so I got this: 6 = 2 1/2 12 = 5 18 = 7.5 24 = 10 The answer = The container would hold 24 liters of juice ?

OpenStudy (apoorvk):

24 in case of 10 times. so one container would contain 24/10, that is 2.4 litres of juice.

OpenStudy (anonymous):

Got it! Thank you!

OpenStudy (apoorvk):

No worries.

OpenStudy (anonymous):

The answer was 5/12.

OpenStudy (apoorvk):

yeah, i told you that itself --> 2.5/6 = 5/12 --> what were you thinking all the time??

OpenStudy (anonymous):

Sorry I meant the correct answer was either 12/5 or 2 2/5